OpenStudy (anonymous):

what is the quotient of \[\frac{ \frac{ x^{2}-1}{ x^{2}+1} }{ \frac{ x^{2}-2x+1 }{ x^{2}+2x+1 } }\]

OpenStudy (anonymous):

\[=\frac{ (x-1)(x+1) }{ x^2+1 } \div \frac{ (x-1)(x-1) }{(x+1)(x+1) }=\frac{ \cancel{(x-1)}(x+1) }{ x^2+1 } \times \frac{(x+1)(x+1) }{\cancel{(x-1)}(x-1) }\]\[=\frac{ (x+1)^3 }{ (x^2+1)(x-1) }\] @jake.rachlin

OpenStudy (anonymous):

you are a lifesaver! @genius12

OpenStudy (anonymous):

Medal/fan/testimonial? =[ @jake.rachlin

OpenStudy (anonymous):

i did that thank you so much @genius12

OpenStudy (anonymous):

Np. Btw, you do understand what I typed up right? Like do you understand the steps? @jake.rachlin

OpenStudy (anonymous):

yes 100% better than my teacher thanks!!!! @genius12